# Unlearning Inversion Attacks for Graph Neural Networks

### v1

@@ 3 Preliminaries
We focus on the edge unlearning setting wu2023certified , where $\Delta\mathcal{V}=\emptyset$, i.e., only edges are removed. This setting captures practical scenarios such as users requesting the removal of private friendship links from social media or the deletion of sensitive purchase records from recommender systems.
4 Problem Formulation
4.1 Threat Model
@@
Attacker’s Knowledge and Capability. We consider a black-box setting, motivated by the widespread deployment of machine learning models as a service via APIs shokri2017membership ; truex2019demystifying . The attacker can query the unlearned GNN to obtain output probabilities $f_{\mathcal{G}_{\mathrm{un}}}(\cdot;\boldsymbol{\theta}_{\mathrm{un}})$ for target nodes in $\mathcal{V}$. Additionally, the attacker has partial access to the unlearned graph $\mathcal{G}_{\mathrm{un}}$. For any pair of target nodes $v_{i},v_{j}\in\mathcal{V}$, the attacker has access to the following information: (i) The model output probabilities $f_{\mathcal{G}_{\mathrm{un}}}(v_{i};\boldsymbol{\theta}_{\mathrm{un}})$ and $f_{\mathcal{G}_{\mathrm{un}}}(v_{j};\boldsymbol{\theta}_{\mathrm{un}})$; (ii) The input features $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$; and (iii) A subset of the $k$-hop neighborhood of $v_{i}$ and $v_{j}$, i.e., $\widehat{\mathcal{N}}^{(k)}(v_{i})$ and $\widehat{\mathcal{N}}^{(k)}(v_{j})$. Despite extending the typical probability-only black-box attack setting in (i), this level of access is realistic in practice, since (ii) can be obtained from users’ public profiles on online social platforms, and (iii) can be retrieved by querying public friend lists or social connections. Furthermore, due to the wide availability of public social network datasets dai2020ginger ; aliapoulios2021large ; nielsen2022mumin , we assume that the attacker also has access to a shadow dataset $\mathcal{G}^{\mathrm{sha}}$ with a distribution similar to $\mathcal{G}_{\mathrm{orig}}$. This shadow graph can be used to train a surrogate model to support the unlearning inversion attack.
4.2 Graph Unlearning Inversion
@@
The query set is a set of $Q$ node pairs of interest, defined as $\mathcal{Q}:=\{(v_{i_{k}},v_{j_{k}}):v_{i_{k}},v_{j_{k}}\in\mathcal{V}\}_{k=1}^{Q}$.
@@
Let $f_{\mathcal{G}_{\mathrm{un}}}(\cdot;\boldsymbol{\theta}_{\mathrm{un}})$ be an unlearned GNN model where the parameters $\boldsymbol{\theta}_{\mathrm{un}}$ are obtained from Eq. (2), and $\Delta\mathcal{E}$ denotes the unlearned edge set. For a given query set $\mathcal{Q}$, the adversary’s partial knowledge on the unlearned graph $\mathcal{G}_{\mathrm{un}}$ is defined as a tuple $\mathcal{K}_{\mathcal{Q}}:=(\mathcal{P}_{\mathcal{Q}},\mathcal{F}_{\mathcal{Q}})$, where: (i) Probability set $\mathcal{P}_{\mathcal{Q}}:=\{f_{\mathcal{G}_{\mathrm{un}}}(v_{i};\boldsymbol{\theta}_{\mathrm{un}}):v_{i}\in\widehat{\mathcal{N}}^{(k)}(v_{j}),v_{j}\in\mathcal{Q}\}$ represents the model output probabilities for nodes in the $k$-hop neighborhood of the query nodes; and (ii) Feature set $\mathcal{F}_{\mathcal{Q}}:=\{\mathbf{x}_{i}:v_{i}\in\mathcal{Q}\}$ contains the input features of nodes in the query pairs.
With the above definition, the graph unlearning inversion problem is defined as
@@
Let $\mathcal{Q}$ be the adversary’s node pairs of interest (Definition 4.1), with partial knowledge $\mathcal{K}_{\mathcal{Q}}$ on the unlearned graph $\mathcal{G}_{\mathrm{un}}$ (Definition 4.2). Suppose the adversary also has access to a shadow graph $\mathcal{G}^{\mathrm{sha}}$ drawn from a distribution similar to $\mathcal{G}_{\mathrm{orig}}$. The goal of the graph unlearning inversion attack is to predict, for each $(v_{i},v_{j})\in\mathcal{Q}$, whether $(v_{i},v_{j})\in\mathcal{E}_{\mathrm{orig}}$ (label $1$) or $(v_{i},v_{j})\notin\mathcal{E}_{\mathrm{orig}}$ (label 00).
@@
More specifically, the query set $\mathcal{Q}$ can be divided into three disjoint subsets: (i) $\mathcal{Q}^{+}_{\mathrm{un}}:=\{(v_{i},v_{j}):(v_{i},v_{j})\in\Delta\mathcal{E}\}$, representing the positive unlearned edges; (ii) $\mathcal{Q}^{+}_{\mathrm{me}}:=\{(v_{i},v_{j}):(v_{i},v_{j})\in\mathcal{E}_{\mathrm{orig}}\setminus\Delta\mathcal{E}\}$, representing the remaining membership edges in the original graph; and (iii) $\mathcal{Q}^{-}:=\{(v_{i},v_{j}):(v_{i},v_{j})\notin\mathcal{E}_{\mathrm{orig}}\}$, denoting non-member (negative) pairs. The goal of this work is to distinguish both $\mathcal{Q}^{+}_{\mathrm{un}}$ and $\mathcal{Q}^{+}_{\mathrm{me}}$ from $\mathcal{Q}^{-}$, enabling accurate membership inference on both positive subsets. This highlights a key difference between our work and prior MIA methods for GNNs he2021node ; he2021stealing ; zhang2023demystifying , which focus solely on distinguishing $\mathcal{Q}^{+}_{\mathrm{me}}$ from $\mathcal{Q}^{-}$ and may fall short of inferring $\mathcal{Q}^{+}_{\mathrm{un}}$.
5 Proposed Method
Figure 2: Illustration of the proposed TrendAttack.
@@ 5 Proposed Method
In this section, we first present the key motivation behind our method, focusing on the adaptive threshold and confidence pitfalls. We then introduce our proposed TrendAttack framework. An illustration of design motivation and attack process of TrendAttack is in Figure 2.
5.1 Design Motivation
As discussed in Section 4.2, a primary challenge that differentiates graph unlearning inversion attacks from traditional MIA on GNNs is the need to distinguish both unlearned edges in $\mathcal{Q}_{\mathrm{un}}^{+}$ and other membership edges in $\mathcal{Q}_{\mathrm{mem}}^{+}$ from non-member node pairs in $\mathcal{Q}^{-}$. To address this, we identify two key observations: one explaining why existing MIA methods fall short, and another guiding how to improve upon them using insights from confidence trends.
Probability Similarity Gap. Existing MIA methods infer the presence of a link between $v_{i}$ and $v_{j}$ by evaluating the similarity between their input features $\mathrm{sim}(\mathbf{x}_{i},\mathbf{x}_{j})$ he2021stealing and/or the similarity between their predicted probabilities $\mathrm{sim}(\mathbf{p}_{i},\mathbf{p}_{j})$ he2021stealing ; olatunji2021membership , where $\mathbf{p}_{i}$ is shorthand for the GNN output $f_{\mathcal{G}_{\mathrm{un}}}(v_{i};\boldsymbol{\theta}_{\mathrm{un}})$. In the link unlearning setting, since input features remain unchanged after removing $\Delta\mathcal{E}$ from $\mathcal{G}_{\mathrm{orig}}$ and deriving $\boldsymbol{\theta}_{\mathrm{un}}$ from $\boldsymbol{\theta}_{\mathrm{orig}}$, feature similarity $\mathrm{sim}(\mathbf{x}_{i},\mathbf{x}_{j})$ provides a consistent signal for identifying both unlearned and other membership edges. This suggests that existing feature-based MIA techniques remain applicable.
However, it is essential to carefully examine how probability similarity $\mathrm{sim}(\mathbf{p}_{i},\mathbf{p}_{j})$ behaves across different edge types. For a removed edge $(v_{i},v_{j})\in\mathcal{Q}^{+}_{\mathrm{un}}$, which is excluded from the training graph (i.e., $(v_{i},v_{j})\notin\mathcal{G}_{\mathrm{un}}$) and unlearned from the parameters $\boldsymbol{\theta}_{\mathrm{un}}$, the link’s influence on model predictions tends to be weaker than that of other membership edges in $\mathcal{Q}^{+}_{\mathrm{mem}}$. Due to approximation errors that are inherent in practical unlearning techniques chien2022certified ; wu2023gif , some residual signal from $(v_{i},v_{j})$ may still persist in the model’s predictions. These observations lead to the following hypothesis:
@@ 5.1 Design Motivation
$\mathrm{ProbSim}(\mathcal{Q}^{-})<\mathrm{ProbSim}(\mathcal{Q}^{+}_{\mathrm{un}})<\mathrm{ProbSim}(\mathcal{Q}^{+}_{\mathrm{mem}}),$
@@
This claim is empirically verified by our preliminary study in Appendix E.
Confidence Pitfall. Our analysis of the probability similarity gap (Claim 5.1) suggests that unlearned edges $\mathcal{Q}^{+}_{\mathrm{un}}$ and remaining membership edges $\mathcal{Q}^{+}_{\mathrm{mem}}$ require different probability-similarity thresholds for accurate inference. The challenge, however, is that the attacker cannot directly tell whether a given query $(v_{i},v_{j})$ belongs to $\mathcal{Q}^{+}_{\mathrm{un}}$ or $\mathcal{Q}^{+}_{\mathrm{mem}}$. To resolve this problem, we present the following intuition:
@@
The average model confidence of nodes appearing in unlearned edges is lower than that of other nodes, i.e., $\mathrm{AvgConf}(\mathcal{V}_{\mathcal{Q}^{+}_{\mathrm{un}}})<\mathrm{AvgConf}(\mathcal{V}\setminus\mathcal{V}_{\mathcal{Q}^{+}_{\mathrm{un}}})$, where $\mathcal{V}_{\mathcal{Q}_{0}}:=\{v_{i}:(v_{i},v_{j})\in\mathcal{Q}_{0}~{}\mathrm{or}~{}(v_{j},v_{i})\in\mathcal{Q}_{0}\}$ is the set of all nodes involved in query subset $\mathcal{Q}_{0}$, and $\mathrm{AvgConf}(\mathcal{V}_{0}):=\frac{1}{|\mathcal{V}_{0}|}\sum_{v_{i}\in\mathcal{V}_{0}}\max_{\ell}\mathbf{p}_{i,\ell}$ is the average model confidence within a node set $\mathcal{V}_{0}$.
For this claim, we provide a detailed theoretical analysis in Appendix C, and then show the empirical evidence and preliminary experiments supporting the claim in Appendix E.
5.2 The Proposed TrendAttack
Based on the two key design principles established in Claim 5.1 and Claim 5.2 for enabling graph unlearning inversion, we now instantiate these principles through concrete model components, resulting in a simple, flexible, and effective inversion framework, named TrendAttack.
Shadow Victim Model Training. To construct an attack model $g(v_{i},v_{j})$ that accurately predicts membership information given partial knowledge of the unlearned graph $\mathcal{K}_{\mathcal{Q}}=(\mathcal{P}_{\mathcal{Q}},\mathcal{F}_{\mathcal{Q}})$, we first simulate the victim model’s behavior using a shadow dataset $\mathcal{G}^{\mathrm{sha}}$ and a shadow victim GNN $f^{\prime}_{\mathcal{G}_{\mathrm{un}}^{\mathrm{sha}}}(\cdot;\boldsymbol{\theta}_{\mathrm{un}}^{\prime})$. This model is trained on a node classification task and then unlearns a small subset of edges $\Delta\mathcal{G}^{\mathrm{sha}}$. Specifically, the pre- and post-unlearning parameters are obtained similarly via Eq. (1) and Eq (2) in Section 3.
This training setup explicitly models unlearning behavior, in contrast to prior MIA approaches that rely solely on the original model $f^{\prime}_{\mathcal{G}^{\mathrm{sha}}_{\mathrm{orig}}}(\cdot;\boldsymbol{\theta}_{\mathrm{orig}}^{\prime})$ and do not account for the effects of unlearning. Using this shadow victim model, we construct a shadow query set $\mathcal{Q}_{\mathrm{sha}}$ and its associated partial knowledge $\mathcal{K}_{\mathcal{Q}_{\mathrm{sha}}}=(\mathcal{P}_{\mathcal{Q}_{\mathrm{sha}}},\mathcal{F}_{\mathcal{Q}_{\mathrm{sha}}})$ from the features, connectivity, and outputs of $f^{\prime}$ on $\mathcal{G}^{\mathrm{sha}}$, which are then used to train the attack model.
Attack Model. It is well established that membership between $v_{i}$ and $v_{j}$ can be inferred from the similarity between their features and output probabilities olatunji2021membership ; he2021stealing . Since prior MIA methods have developed a variety of similarity computation frameworks for this task, we adopt a general formulation that computes a scalar similarity score between $v_{i}$ and $v_{j}$ as $\phi([\mathbf{x}_{i}~{}||~{}\mathbf{p}_{i}],[\mathbf{x}_{j}~{}||~{}\mathbf{p}_{j}])$.

### v2

@@ 3. Preliminaries
where the right optimization problem denotes retrain from scratch on the unlearned graph. We focus on the edge unlearning setting (wu2023certified), where $\Delta\mathcal{V}=\emptyset$, i.e., only edges are removed. This setting captures practical scenarios such as users requesting the removal of private friendship links from social media or the deletion of sensitive purchase records from recommender systems.
4. Problem Formulation
In this section, we begin by describing the threat model associated with the link-level graph unlearning inversion attack, and then present a formal problem definition.
4.1. Threat Model
@@
Attacker’s Knowledge and Capability. We consider a black-box setting, motivated by the widespread deployment of machine learning models as a service via APIs (shokri2017membership; truex2019demystifying). The attacker can query the unlearned GNN to obtain output probabilities $f_{\mathcal{G}_{\mathrm{un}}}(\cdot;\bm{\theta}_{\mathrm{un}})$ for target nodes in $\mathcal{V}$. Additionally, the attacker has partial access to the unlearned graph $\mathcal{G}_{\mathrm{un}}$. For any pair of target nodes $v_{i},v_{j}\in\mathcal{V}$, the attacker has access to the following information: (i) The model output probabilities $f_{\mathcal{G}_{\mathrm{un}}}(v_{i};\bm{\theta}_{\mathrm{un}})$ and $f_{\mathcal{G}_{\mathrm{un}}}(v_{j};\bm{\theta}_{\mathrm{un}})$; (ii) The input features $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$; and (iii) A subset of the $k$-hop neighborhood of $v_{i}$ and $v_{j}$, i.e., $\widehat{\mathcal{N}}^{(k)}(v_{i})$ and $\widehat{\mathcal{N}}^{(k)}(v_{j})$. Furthermore, similar to many previous privacy attacks on GNNs (he2021stealing; olatunji2021membership; dai2023unified) that use a shadow dataset to train a surrogate model, we also assume (iv) the attacker has access to a shadow dataset $\mathcal{G}^{\mathrm{sha}}$ with a distribution similar to the training graph $\mathcal{G}_{\mathrm{orig}}$ before unlearning.
Discussion. Although our attack setting extends the typical probab
-ility-only black-box attack in (i), this level of attacker knowledge is realistic in practice, since (ii) can be obtained from users’ public profiles on social media platforms, and (iii) can be retrieved by querying public friend lists or social connections.
We note that such shadow datasets in (iv) are often accessible in practice. For example, in social network attacks, many public datasets are available (dai2020ginger; aliapoulios2021large; nielsen2022mumin), enabling us to train an attack model on one platform and apply it to another (e.g., from X (Twitter) to Facebook), or transfer it across regions (e.g., Facebook networks in the US to those in the UK). In FinTech scenarios (e.g., GNN-based APIs for credit recommendation or fraud detection), attackers could build shadow datasets from public transaction networks (e.g., blockchain data) (weber2019anti; pareja2020evolvegcn; chai2023towards), train TrendAttack models on surrogate victim models, and transfer them to real victim models to infer sensitive transactions from API outputs. For instance, the Bitcoin transaction network can serve as a shadow dataset, allowing the attack model to be transferred to other De-Fi ecosystems such as Ethereum or newly emerging platforms.
4.2. Graph Unlearning Inversion
@@
The query set is a set of $Q$ node pairs of the adversary’s interest, defined as $\mathcal{Q}:=\{(v_{i_{k}},v_{j_{k}}):v_{i_{k}},v_{j_{k}}\in\mathcal{V}\}_{k=1}^{Q}$, where $|\mathcal{Q}|=Q$.
@@
Let $f_{\mathcal{G}_{\mathrm{un}}}(\cdot;\bm{\theta}_{\mathrm{un}})$ be an unlearned GNN model where the parameters $\bm{\theta}_{\mathrm{un}}$ are obtained from Eq. (3), and $\Delta\mathcal{E}$ denotes the unlearned edge set. For a given query set $\mathcal{Q}$, the adversary’s partial knowledge on the unlearned graph $\mathcal{G}_{\mathrm{un}}$ is defined as a tuple $\mathcal{K}_{\mathcal{Q}}:=(\mathcal{P}_{\mathcal{Q}},\mathcal{F}_{\mathcal{Q}})$, where: (i) Probability set $\mathcal{P}_{\mathcal{Q}}:=\{f_{\mathcal{G}_{\mathrm{un}}}(v_{i};\bm{\theta}_{\mathrm{un}}):v_{i}\in\widehat{\mathcal{N}}^{(k)}(v_{j}),v_{j}\in\mathcal{Q}\}$ represents the black-box model output probabilities for nodes in the $k$-hop neighborhood of the query nodes; and (ii) Feature set $\mathcal{F}_{\mathcal{Q}}:=\{\mathbf{x}_{i}:v_{i}\in\mathcal{Q}\}$ contains the input features of nodes in the query pairs.
In this setting, the attacker aims to find out whether node pairs in the query set $\mathcal{Q}$ (Definition 4.1) were connected in the original graph $\mathcal{G}_{\mathrm{orig}}$ before edge removal and GNN unlearning. Under the partial knowledge setting in Definition 4.2, the adversary has only black-box access to the unlearned model $f(\cdot;\bm{\theta}_{\mathrm{un}})$ on the clean graph $\mathcal{G}_{\mathrm{un}}$ (i.e., the graph without the unlearned edges), which contains minimal information about removed edges, making our attack highly non-trivial. Compared with the strictest MIA setting (i.e., where only black-box model outputs are available) (olatunji2021membership), the attacker also knows the IDs of part of neighboring nodes in the query set $\mathcal{Q}$ for black-box model queries, and has access to the input features of the nodes. Note that such assumptions are common in membership inference attacks, e.g., StealLink (he2021stealing) also assumes access to node features and a partial view of the attack graph.
With the above definitions, the graph unlearning inversion problem is then formalized as:
@@
Let $\mathcal{Q}$ be the adversary’s node pairs of interest (Definition 4.1), with partial knowledge $\mathcal{K}_{\mathcal{Q}}$ on the unlearned graph $\mathcal{G}_{\mathrm{un}}$ (Definition 4.2). Suppose the adversary also has access to a shadow graph $\mathcal{G}^{\mathrm{sha}}$ drawn from a distribution similar to $\mathcal{G}_{\mathrm{orig}}$. The goal of the graph unlearning inversion attack is to predict, for each $(v_{i},v_{j})\in\mathcal{Q}$, whether $(v_{i},v_{j})\in\mathcal{E}_{\mathrm{orig}}$ (label $1$) or $(v_{i},v_{j})\notin\mathcal{E}_{\mathrm{orig}}$ (label 0).
@@
More specifically, the query set $\mathcal{Q}$ can be divided into three disjoint subsets: (i) $\mathcal{Q}^{+}_{\mathrm{un}}:=\{(v_{i},v_{j}):(v_{i},v_{j})\in\Delta\mathcal{E}\}$, representing the positive unlearned edges; (ii) $\mathcal{Q}^{+}_{\mathrm{mem}}:=\{(v_{i},v_{j}):(v_{i},v_{j})\in\mathcal{E}_{\mathrm{orig}}\setminus\Delta\mathcal{E}\}$, representing the remaining membership edges in the original graph; and (iii) $\mathcal{Q}^{-}:=\{(v_{i},v_{j}):(v_{i},v_{j})\notin\mathcal{E}_{\mathrm{orig}}\}$, denoting non-member (negative) pairs. The goal of this work is to distinguish both $\mathcal{Q}^{+}_{\mathrm{un}}$ and $\mathcal{Q}^{+}_{\mathrm{mem}}$ from $\mathcal{Q}^{-}$, enabling accurate membership inference on both positive subsets. This highlights a key difference between our work and prior MIA methods for GNNs (he2021node; he2021stealing; zhang2023demystifying), which focus solely on distinguishing $\mathcal{Q}^{+}_{\mathrm{mem}}$ from $\mathcal{Q}^{-}$ and may fall short of inferring $\mathcal{Q}^{+}_{\mathrm{un}}$. Moreover, our work also differs substantially from previous unlearning inversion attacks (bertran2024reconstruction; hu2024learn). A summary is provided in the Remark 4.4 below, and detailed comparisons on attack settings can be found in Supplementary H.1.
Remark 4.4 (Strictness of Attack Setting).
Despite using a shadow dataset $\mathcal{G}^{\mathrm{sha}}$, our attack still works under one of the strictest settings in the unlearning inversion literature. Unlike previous works (bertran2024reconstruction; hu2024learn) that require white-box or black-box access to both the original model $f(\cdot;\bm{\theta}_{\mathrm{orig}})$ and the unlearned model $f(\cdot;\bm{\theta}_{\mathrm{un}})$, we only require black-box access to the unlearned model $f(\cdot;\bm{\theta}_{\mathrm{orig}})$.
Remark 4.5 (Difference to GNN MIAs).
Existing GNN MIA methods (he2021node; he2021stealing; zhang2023demystifying) aim to tell whether $(v_{i},v_{j})\in\mathcal{E}_{\mathrm{orig}}$ by using the black-box output of the original GNN model $f(\cdot;\bm{\theta}_{\mathrm{orig}})$, which may memorize such edges. In contrast, our attack in Definition 4.3 works in a harder setting: we infer whether $(v_{i},v_{j})\in\mathcal{E}_{\mathrm{orig}}$ from the unlearned model $f(\cdot;\bm{\theta}_{\mathrm{un}})$, with minimal knowledge about the removed edges $\Delta\mathcal{E}\subseteq\mathcal{E}_{\mathrm{orig}}$. Remarkably, even if $(v_{i},v_{j})\in\Delta\mathcal{E}$ and has been unlearned, we can still recover it accurately.
5. Proposed Method
Figure 2. Illustration of the proposed TrendAttack.
@@ 5. Proposed Method
In this section, we first present the key motivation behind our method, focusing on the adaptive threshold and confidence pitfalls. We then introduce our proposed TrendAttack framework. An illustration of the proposed TrendAttack is in Figure 2.
5.1. Design Motivation
As discussed in Section 4.2, a key challenge that differentiates graph unlearning inversion attacks from traditional MIAs on GNNs is the need to distinguish both unlearned edges in $\mathcal{Q}_{\mathrm{un}}^{+}$ and memorized edges in $\mathcal{Q}_{\mathrm{mem}}^{+}$ from non-edge pairs in $\mathcal{Q}^{-}$. However, these two types of positive edges exhibit different behaviors: unlearned edges typically require a lower similarity threshold to be separated from $\mathcal{Q}^{-}$, while memorized edges require a higher threshold. This raises the question of how to adaptively decide which threshold to apply.
Our first claim confirms the necessity of different thresholds, and our second claim shows that confidence trends provide a useful signal for classifying whether an edge belongs to $\mathcal{Q}_{\mathrm{un}}^{+}$ or $\mathcal{Q}_{\mathrm{mem}}^{+}$, thereby guiding the choice of threshold.
Probability Similarity Gap. Existing MIA methods typically decide whether a link exists between $v_{i}$ and $v_{j}$ by measuring the similarity of their input features $\mathrm{sim}(\mathbf{x}_{i},\mathbf{x}_{j})$ (he2021stealing) and/or the similarity of their output probabilities $\mathrm{sim}(\mathbf{p}_{i},\mathbf{p}_{j})$ (he2021stealing; olatunji2021membership), where $\mathbf{p}_{i}=f_{\mathcal{G}_{\mathrm{un}}}(v_{i};\bm{\theta}_{\mathrm{un}})$ denotes the GNN output. While feature similarity $\mathrm{sim}(\mathbf{x}_{i},\mathbf{x}_{j})$ is not affected by unlearning and can still be applied in our setting, it does not fully capture graph structure information. Probability similarity $\mathrm{sim}(\mathbf{p}_{i},\mathbf{p}_{j})$, on the other hand, encodes both node features and the structural signals learned by the GNN, and thus plays a central role in distinguishing edges from non-edges.
However, unlearning would affect the similarity of probability. When an edge $(v_{i},v_{j})$ belongs to $\Delta\mathcal{E}$ and is removed from the graph, its structural contribution to the model is explicitly erased during unlearning. As a result, the similarity between $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ weakens compared to edges that remain memorized by the GNN in $\mathcal{Q}_{\mathrm{mem}}^{+}$. These considerations lead us to the following claim, which is empirically verified by our preliminary study in Supplementary E:
@@ 5.1. Design Motivation
[…]
@@
The intuition for this claim is that, memorized edges $\mathcal{Q}^{+}_{\mathrm{mem}}$ tend to show the highest probability similarity because the model has fully maintained their structural signals. After unlearning, the edges in $\mathcal{Q}^{+}_{\mathrm{un}}$ lose part of this signal, so their similarity becomes lower than memorized edges. However, practical unlearning methods are often approximate (chien2022certified; wu2023gif), which means that a small residual effect of the removed edges may still remain in the model, preventing their similarity from dropping to the level of non-existent edges $\mathcal{Q}^{-}$. In contrast, non-edges naturally exhibit the lowest similarity because they lack any structural signal.
This gap indicates that, to effectively perform the privacy attack with model output probabilities, one must first distinguish between unlearned and memorized edges, and then apply an adaptive similarity threshold when predicting edges versus non-edges.
Confidence Pitfall. Our analysis of the probability similarity gap (Claim 5.1) suggests that unlearned edges $\mathcal{Q}^{+}_{\mathrm{un}}$ and remaining memorized edges $\mathcal{Q}^{+}_{\mathrm{mem}}$ require different probability-similarity thresholds for accurate prediction. The remaining challenge, however, is that the attacker cannot directly tell whether a given query $(v_{i},v_{j})$ belongs to $\mathcal{Q}^{+}_{\mathrm{un}}$ or $\mathcal{Q}^{+}_{\mathrm{mem}}$, which makes it difficult to decide which similarity threshold to choose for the query $(v_{i},v_{j})$.
To address this, we analyze how removing a specific edge $(v_{i},v_{j})$ affects model outputs, using the widely used analytical framework of influence functions (koh2017understanding; wu2023gif; wu2023certified). This analysis separates two effects: (i) the immediate impact of dropping the edge from the training graph $\mathcal{G}_{\mathrm{orig}}$, and (ii) the subsequent adjustment of model parameters by the unlearning procedure. We focus on a linear GCN model, which, despite its simplicity, captures the core behavior of many GNN architectures by choosing a different propagation matrix $\mathbf{C}$ (see Remark D.2 in Supplementary D).
Theorem 5.2 (Single‐Edge to Single‐Output Influence, Informal).
Let $f(\mathbf{C},\mathbf{X};\mathbf{w}^{\star}):=\mathbf{C}\mathbf{X}\mathbf{w}^{\star}$ be a linear GCN with propagation matrix $\mathbf{C}\in\mathbb{R}^{n\times n}$ and parameters $\mathbf{w}^{\star}$ obtained by least‐squares on labels $\mathbf{y}\in\mathbb{R}^{d}$. The influence of an undirected edge $(v_{i},v_{j})\in\Delta\mathcal{E}$ on the model output $\mathbf{p}_{k}:=f(\mathbf{C},\mathbf{X};\mathbf{w}^{\star})_{k}$ of node $v_{k}\in\mathcal{V}$ can be decomposed as follows:
$\mathcal{I}(\mathbf{p}_{k})=$ $~\underbrace{\bm{1}{\{v_{k}=v_{i}\}}\cdot(\mathbf{x}^{\top}_{i}\mathbf{w}^{\star})+\bm{1}{\{v_{k}=v_{j}\}}\cdot(\mathbf{x}^{\top}_{j}\mathbf{w}^{\star})}_{\mathrm{edge~influence}}$
$~~~~-\underbrace{\langle(\mathbf{x}_{j}^{\top}\mathbf{w}^{\star})\mathbf{z}_{i}+(\mathbf{x}_{i}^{\top}\mathbf{w}^{\star})\mathbf{z}_{j},\mathbf{z}_{k}\rangle_{\mathbf{H}^{-1}}}_{\mathrm{magnitude~weight~influence}}$
$~~~~+\underbrace{\langle(y_{j}-\mathbf{z}_{j}^{\top}\mathbf{w}^{\star})\mathbf{x}_{i}+(y_{i}-\mathbf{z}_{i}^{\top}\mathbf{w}^{\star})\mathbf{x}_{j},\mathbf{z}_{k}\rangle_{\mathbf{H}^{-1}}}_{\mathrm{error~weight~influence}},$
where $\mathbf{Z}:=\mathbf{C}\mathbf{X}$ and $\mathbf{z}_{l}$ is the ll‐th row of $\mathbf{Z}$, $\bm{1}{\{\cdot\}}$ denotes the indicator function, $\mathbf{H}$ is the Hessian of the least squares loss evaluated at $\mathbf{w}^{\star}$, and $\langle\cdot,\cdot\rangle_{\mathbf{A}}$ denotes the weighted inner product with any positive semi-definite (PSD) matrix $\mathbf{A}$.
Proof.
Please see Theorem D.11 in Supplementary D. ∎
In the theorem above, the third error weight influence term is governed by the empirical residuals at the optimal weight, $(y_{i}-\mathbf{z}_{i}^{\top}\mathbf{w}^{\star})$ and $(y_{j}-\mathbf{z}_{j}^{\top}\mathbf{w}^{\star})$. Under the assumption of a well‐trained (i.e., learnable) graph ML problem, these residuals become negligibly small, making the error weight influence term nearly zero.
Meanwhile, the first edge influence term activates only when $v_{k}$ coincides with one of the endpoints $v_{i}$ or $v_{j}$, producing a direct and large perturbation. The second magnitude weight influence term, which depends on the inner‐product similarity $\langle(\mathbf{x}_{j}^{\top}\mathbf{w}^{\star})\mathbf{z}_{i}+(\mathbf{x}_{i}^{\top}\mathbf{w}^{\star})\mathbf{z}_{j},\mathbf{z}_{k}\rangle_{\mathbf{H}^{-1}}$, is also relatively larger when $v_{k}\in\{v_{i},v_{j}\}$ and smaller otherwise. As a result, for $v_{k}=v_{i}$ or $v_{j}$, the sum of a substantial positive edge influence and a significant negative magnitude weight influence yields a dramatic net effect on the model output. Therefore, the influence on the endpoints of unlearned edges may be more significant than on other nodes. With this, we reasonably assume their confidence will drop and make the following claim:
@@
where $\mathcal{V}_{\mathcal{Q}_{0}}:=\{v_{i}:(v_{i},v_{j})\in\mathcal{Q}_{0}~\mathrm{or}~(v_{j},v_{i})\in\mathcal{Q}_{0}\}$ is the set of all nodes involved in query subset $\mathcal{Q}_{0}$, and $\mathrm{AvgConf}(\mathcal{V}_{0}):=|\mathcal{V}_{0}|^{-1}\sum_{v_{i}\in\mathcal{V}_{0}}\max_{\ell}\mathbf{p}_{i,\ell}$ is the average model confidence of $\mathcal{V}_{0}$.
We also support this claim with a preliminary study that analyzes GNN model confidence at varying distances from unlearned edges on real-world datasets, as detailed in Supplementary E.
5.2. The Proposed TrendAttack
In the previous subsection, we presented two key claims that guide our model design. First, Claim 5.1 indicates that predicting the existence of a link for a node pair $(v_{i},v_{j})$ may require different similarity thresholds for the model outputs: a lower threshold for unlearned edges $(v_{i},v_{j})\in\mathcal{Q}^{+}_{\mathrm{un}}$ and a higher threshold for memorized edges $(v_{i},v_{j})\in\mathcal{Q}^{+}_{\mathrm{mem}}$ to distinguish them from non-existent edges. Second, Claim 5.3 shows that we can distinguish these two types of edges using the model’s confidence trend, which informs which threshold to apply. In this subsection, we instantiate these principles into concrete model components, resulting in a simple, flexible, and effective inversion framework called TrendAttack.
5.2.1. Attack Model